OpenStudy (anonymous):

A rectangular water tank has length 25 feet, wide 10 ft, and depth 10 ft. If the tank is full, how much work does it take to pump all the water out? HELP!! Calculus Problem

OpenStudy (anonymous):

do yu have density?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

how should i approach this

OpenStudy (anonymous):

we know that the area of the base is: 25*10 = 250 If we take a slice out of the tank of thickness dx we will have a slice volume of: Area*height = 250dx water density is roughly 62 Ib/ft^3. So we know that the mass of the slice is: (62)*(250dx) = 15,500dx if he wants to pump all the water out that means from 10ft to 0: \[\int\limits_{10}^{0} 15,500 dx = -775,000 ftlb\] I think that's how it's done, does it lok right?